# Marginal phase transition in decorated single-chain Ising models

### v1

@@ Abstract
The $T-J_{2}$ phase diagrams from the density plot of $\langle\sigma_{i}\rangle$ for the previously studied model (i.e., $J_{2}/J_{1}=1.95$ and $J=0$)  015_Strecka_APPA_20_Ising-diamond-chain ; 016_Strecka_book_chapter is presented in Fig. 3a. The sharp phase crossover from $\langle\sigma_{i}\rangle=-1$ to $+1$ occurs only near the small region with strong geometric frustration ($J_{2}/J_{1}=2$ is the FRI-FRU phase boundary; cf. Fig. 2). The resultant $T_{0}$ is very low ($k_{\mathrm{B}}T_{0}S^{2}/J_{1}\approx 0.036$ for $J_{2}/J_{1}=1.95$). As $J_{2}/J_{1}$ decreases, $T_{0}$ increases ($k_{\mathrm{B}}T_{0}S^{2}/J_{1}\approx 0.821$ for $J_{2}=0$) while the density is spread over a wider temperature interval, meaning larger $\delta T$. However, for the present new model with $J\neq 0$, Fig. 3b shows all the broad crossovers have been turned to be ultra-narrow MPT by increase $J$. The same is observed in the $T-h$ phase diagrams presented in Figs. 4a and 4b for $J_{2}/J_{1}=1.95$ and Figs. 4c and 4d for $J_{2}/J_{1}=0$.
@@
As shown in Fig. 6, both cases resemble a genuine first-order phase transition with the entropy jump and gigantic susceptibility at $T_{0}$, but the entropy per unit cell is flattened at $\ln 2$ and $2\ln 2$ for $J_{2}/J_{1}=1.95$ (Fig. 6a) and $J_{2}=0$ (Fig. 6b), respectively. The former is expected since the case with $J_{2}/J_{1}=1.95$ is located near the FRI-FRU phase boundary in the ground-state phase diagram (the cyan circle in Fig. 2). The FRU phase has the degeneracy of two per unit cell due to the frustrated decorated spins; the crossover is driven by this entropy gain by $\ln 2$ per unit cell 016_Strecka_book_chapter , while the ordinary spins flip from $\sigma_{i}\approx-1$ to $+1$ still being locked together by large $J_{\mathrm{eff}}/k_{\mathrm{B}}T_{0}$. By sharp contrast, the case with $J_{2}=0$ is far away from any phase boundary and the entropy flattening at $2\ln 2$ is astonishing. This exotic phenomenon originates from a hidden remarkably high degeneracy induced by the magnetic field even in the system without geometric frustration: As shown by the red circle in Fig. 2, the case with $J_{2}=0$ is seated right on the extended line of the FRU-SPP phase boundary (red dashed line). This means that once the system is heated out of the FRI ground state, it will be frustrated in choosing between FRU or SPP, resulting in the effective decoupling of the two decorated spins per unit cell from the lattice—that is the entropy gain of $2\ln 2$. Meanwhile, the ordinary spins flip from $\sigma_{i}\approx-1$ to $+1$ and are still locked together by large $J_{\mathrm{eff}}/k_{\mathrm{B}}T_{0}$. In other words, it is this lockup, “calm,” or the buildup of coherence in the order parameter $\sigma_{i}$ of the ordinary spins that makes the decorated spins fully frustrated in response to the heated atmosphere. This is a distinctly new mechanism for driving the MPT. It is also opposite to the zero-temperature critical point recently emphasized as the “half-ice half-fire” state in ferrimagnet-like systems where the ordinary spins are fully frustrated while the decorated spins are forced to be calm by the critical magnetic field Yin_g . Now the FRI regime of the ground-state phase diagram has two paths toward MPT through either the geometric-frustration-driven FRI-FRU or the hidden-frustration-driven FRU-SPP phase boundary. As demonstrated in the density plot of the entropy right above $T_{0}$ (Fig. 7a) and that of the entropy jump at $T_{0}$ (Fig. 7b), the entropy jump of about $\ln 2$ to $2\ln 2$ takes place in most areas of the FRI regime except for weak $h$ and the largest jump occurs approximately along the hidden frustration line. This difference in the entropy jump together with $T_{0}$ and $\delta T$ could be used to train deep neural networks to predict the model parameters of the decorated 1D Ising model with MPT Yin_PRB_23_ML_t-J .
@@
In summary, a simple and general way by including the ferromagnetic $J$ term is found to not only transform all the previously studied systems—with geometric frustration—in the context of pseudo-transition into the MPT cases, but also unexpectedly expose the hidden field-induced frustration to generate the MPT in decorated Ising chains without geometric frustration. With the discoveries of both the spontaneous MPT in the decorated Ising two-leg ladders (which is DNA-like) and the field-driven MPT in the decorated Ising single chains (which is RNA-like), the foundation of the MPT research direction has been solidly established. Given the prominent roles of the Ising model and frustration in understanding collective phenomena in various physical, biological, economical, and social systems, and the prominent roles of 1D systems in research, education, and technology applications, we anticipate that the present new insights to phase transitions and the dynamical actions of frustration will stimulate further research and development about MPT.
@@ Acknowledgements.
The Method
The central quantity of statistic mechanics is the partition function $Z=\mathrm{Tr}\left(e^{-\beta H}\right)$ where $H$ is the Hamiltonian of the system and $\beta=1/(k_{\mathrm{B}}T)$ with $T$ being the temperature and $k_{\mathrm{B}}$ the Boltzmann constant Mattis_book_08_SMMS . The free energy $f(T)=-k_{\mathrm{B}}T\ln Z$ determines many important thermodynamic properties such as the entropy $S=-\partial f/\partial T$, the specific heat $C_{v}=T\partial S/\partial T$, the magnetization $m=-\partial f/\partial h$, and the magnetic susceptibility $\chi=\partial m/\partial h$. Here the OP is $\langle\sigma_{i}\rangle=-(1/h)\partial f/\partial\mu_{a}$.
The partition function of a 1D Ising model can be obtained exactly by using the transfer matrix method Mattis_book_08_SMMS ; Mattis_book_1985 ; Baxter_book_Ising ; 003_Huang_08_book ; 005_Galisova_PRE_15_double-tetrahedral-chain ; 007_Torrico_PRA_16_Ising-XYZ-diamond-chain ; 009_review_Souza_SSC_18_Ising-XYZ-diamond-chain_double-tetrahedral-chain-spin-electron ; 010_Carvalho_JMMM_18_Ising-XYZ-diamond-chain_quantum-entanglement ; 011_Rojas_BJP_20_Ising-Heisenberg-tetrahedral_diamond ; 013_Rojas_PRE_19_previous_4_models ; 014_Rojas_JPC_20_Ising–Heisenberg_spin-1-double-tetrahedral-chain ; 015_Strecka_APPA_20_Ising-diamond-chain ; 015-7_Canova_CzechoslovakJP_04_Ising–Heisenberg_diamond_chain ; 015-8_Canova_JPC_06_Ising–Heisenberg-spin-S-diamond-chain ; 017_Krokhmalskii_PA_21_3-previous-chains_effective_model ; 016_Strecka_book_chapter ; Yin_MPT ; Yin_icecreamcone ; Hutak_PLA_21_trimer ; Yin_g ; Bell_JPC_74_Ising_ferri_1 and is given by
@@
which becomes non-analytic. The difference between Eq. (11) and Eq. (13) takes place in a region of $(T_{0}-\delta T,T_{0}+\delta T)$, where the crossover width $2\delta T$ can be estimated by $|\Upsilon_{-}|=2e^{-2\beta J}\boxed{+-}$ at $T_{0}\pm\delta T$. Following Ref. , an alternative and consistent way to measure $\delta T$ is to find such an order parameter that has a well-defined value space of $[-1,1]$ with the value 00 meaning $T_{0}$ and its inverse slope at $T_{0}$ meaning $\delta T$. It is the magnetization of the ordinary spins given by
$\langle\sigma_{i}\rangle=-\frac{\partial f}{\partial(h\mu_{a})}=\frac{\Upsilon_{-}}{\sqrt{\Upsilon_{-}^{2}+4e^{-4\beta J}\boxed{+-}^{2}}}.$ (14)
@@
Again, it is clear that the crossover width $2\delta T$ decreases exponentially as $J$ increases for fixed finite $T_{0}$. This order parameter provides an accurate, convenient, and microscopic description of MPT. The use of the MPT order parameter accelerates the finding of MPT.
@@
$\begin{split}\delta T=e^{-\frac{2J}{k_{\mathrm{B}}T_{0}}}\frac{2k_{\mathrm{B}}T_{0}^{2}\boxed{+-}}{-\Upsilon_{+}}\biggl{[}h\mu_{a}&+(2\beta J_{1}+\beta h\mu_{b})\tanh(2\beta J_{1}+\beta h\mu_{b})\\ &+(2\beta J_{1}-\beta h\mu_{b})\tanh(2\beta J_{1}-\beta h\mu_{b})\biggr{]}^{-1}_{\beta=\frac{1}{k_{\mathrm{B}}T_{0}}}.\end{split}$ (27)

### v2

@@ II RESULTS AND DISCUSSIONS
The $T-J_{2}$ phase diagrams from the density plot of $\langle\sigma_{i}\rangle$ for the previously studied model (i.e., $J_{2}/J_{1}=1.95$ and $J=0$)  015_Strecka_APPA_20_Ising-diamond-chain ; 016_Strecka_book_chapter is presented in Fig. 3a. For $J=0$, the sharp phase crossover from $\langle\sigma_{i}\rangle=-1$ to $+1$ occurs only near the small region with strong geometric frustration ($J_{2}/J_{1}=2$ is the FRI-FRU phase boundary; cf. Fig. 2). The resultant $T_{0}$ is very low, e.g., $k_{\mathrm{B}}T_{0}S^{2}/|J_{1}|\approx 0.036$ for $J_{2}/J_{1}=1.95$ [c.f., Equation (3)]. As $J_{2}/J_{1}$ decreases, $T_{0}$ increases ($k_{\mathrm{B}}T_{0}S^{2}/|J_{1}|\approx 0.821$ for $J_{2}=0$) while the density is spread over a wider temperature interval, meaning larger $\delta T$. However, for the present new model with $J>0$, Fig. 3b shows that all the broad crossovers have been turned to be ultranarrow by increasing $J$. The same is observed in the $T-h$ phase diagrams presented in Figs. 4a and 4b for $J_{2}/J_{1}=1.95$ and Figs. 4c and 4d for $J_{2}/J_{1}=0$.
@@
As shown in Fig. 6, both cases resemble a genuine first-order phase transition with the entropy jump and gigantic susceptibility at $T_{0}$, but the entropy per unit cell is flattened at $\ln 2$ and $2\ln 2$ for $J_{2}/J_{1}=1.95$ (Fig. 6a) and $J_{2}=0$ (Fig. 6b), respectively. The former is expected since the case with $J_{2}/J_{1}=1.95$ is located near the FRI-FRU phase boundary in the ground-state phase diagram (the cyan circle in Fig. 2). The FRU phase has the degeneracy of two per unit cell due to the frustrated decorated spins; the crossover is driven by this entropy gain by $\ln 2$ per unit cell 016_Strecka_book_chapter , while the ordinary spins flip from $\sigma_{i}\approx-1$ to $+1$ still being locked together by large $J_{\mathrm{eff}}/k_{\mathrm{B}}T_{0}$. By sharp contrast, the case with $J_{2}=0$ is far away from any phase boundary and the entropy flattening at $2\ln 2$ is astonishing. This exotic phenomenon originates from a hidden remarkably high degeneracy induced by the magnetic field even in the system without geometric frustration: As shown by the red circle in Fig. 2, the case with $J_{2}=0$ is seated right on the extended line of the FRU-SPP phase boundary (red dashed line). This means that once the system is heated out of the FRI ground state, it will be frustrated in choosing between FRU or SPP, resulting in the effective decoupling of the two decorated spins per unit cell from the lattice—that is the entropy gain of $2\ln 2$. Meanwhile, the ordinary spins flip from $\sigma_{i}\approx-1$ to $+1$ and are still locked together by large $J_{\mathrm{eff}}/k_{\mathrm{B}}T_{0}$. In other words, it is this lockup, “calm,” or the buildup of coherence in the order parameter $\sigma_{i}$ of the ordinary spins that makes the decorated spins fully frustrated in response to the heated atmosphere. This is a distinctly new mechanism for driving the CPT. It is also opposite to the zero-temperature critical point recently emphasized as the “half-ice half-fire” state in ferrimagnet-like systems where the ordinary spins are fully frustrated while the decorated spins are forced to be calm by the critical magnetic field Yin_g . Now the FRI regime of the ground-state phase diagram has two paths toward CPT through either the geometric-frustration-driven FRI-FRU or the hidden-frustration-driven FRU-SPP phase boundary. As demonstrated in the density plot of the entropy right above $T_{0}$ (Fig. 7a) and that of the entropy jump at $T_{0}$ (Fig. 7b), the entropy jump of about $\ln 2$ to $2\ln 2$ takes place in most areas of the FRI regime except for weak $h$ and the largest jump occurs approximately along the hidden frustration line. This difference in the entropy jump together with $T_{0}$ and $\delta T$ could be used to train deep neural networks to predict the model parameters of the decorated 1D Ising model with CPT Yin_PRB_23_ML_t-J .
III Summary
@@
In summary, a simple and general way by including the ferromagnetic $J$ term is found to not only transform all the previously studied systems—with geometric frustration—in the context of pseudo-transition into the ultranarrow phase crossover that possesses the highly desirable features: $T_{0}$ and $2\delta T$ can be independently controlled by different parameters and different decoration methods, but also unexpectedly expose the hidden field-induced frustration to generate the ultranarrow phase crossover in decorated Ising chains without the conventional geometric frustration. With the discoveries of both the spontaneous ultranarrow phase crossover in the decorated Ising two-leg ladders (which is DNA-like) and the field-driven ultranarrow phase crossover in the decorated Ising single chains (which is RNA-like), the foundation of the research direction in ultranarrow phase crossover has been solidly established. Given the prominent roles of the Ising model and frustration in understanding collective phenomena in various physical, biological, economical, and social systems, and the prominent roles of 1D systems in research, education, and technology applications, as well as the recent technological advancement that the Ising model has already been implemented in various physical systems Ising_FPGA ; Pierangeli_IsingMachine_PRL19 ; Bernien_Nature_17_Rydberg , we anticipate that the present new insights to phase transitions and frustration effects will stimulate further research and development about ultranarrow phase crossover and its applications.
@@ Acknowledgements.
Appendix A The Method
The central quantity of statistic mechanics is the partition function $Z=\mathrm{Tr}\left(e^{-\beta H}\right)$ where $H$ is the Hamiltonian of the system and $\beta=1/(k_{\mathrm{B}}T)$ with $T$ being the temperature and $k_{\mathrm{B}}$ the Boltzmann constant Mattis_book_08_SMMS . The free energy per unit cell $f(T)=-{\color[rgb]{0,0,0}\frac{1}{N}}k_{\mathrm{B}}T\ln Z$ (where $N$ is the number of unit cells) determines many important thermodynamic properties such as the entropy $S=-\partial f/\partial T$, the specific heat $C_{v}=T\partial S/\partial T$, the magnetization $m=-\partial f/\partial h$, and the magnetic susceptibility $\chi=\partial m/\partial h$. Here the OP is $\langle\sigma_{i}\rangle=-(1/h)\partial f/\partial\mu_{a}$.
The partition function of a 1D Ising model can be obtained exactly by using the transfer matrix method Mattis_book_08_SMMS ; Mattis_book_1985 ; Baxter_book_Ising ; 003_Huang_08_book ; 005_Galisova_PRE_15_double-tetrahedral-chain ; 007_Torrico_PRA_16_Ising-XYZ-diamond-chain ; 009_review_Souza_SSC_18_Ising-XYZ-diamond-chain_double-tetrahedral-chain-spin-electron ; 010_Carvalho_JMMM_18_Ising-XYZ-diamond-chain_quantum-entanglement ; 011_Rojas_BJP_20_Ising-Heisenberg-tetrahedral_diamond ; 013_Rojas_PRE_19_previous_4_models ; 014_Rojas_JPC_20_Ising–Heisenberg_spin-1-double-tetrahedral-chain ; 015_Strecka_APPA_20_Ising-diamond-chain ; 015-7_Canova_CzechoslovakJP_04_Ising–Heisenberg_diamond_chain ; 015-8_Canova_JPC_06_Ising–Heisenberg-spin-S-diamond-chain ; 017_Krokhmalskii_PA_21_3-previous-chains_effective_model ; 016_Strecka_book_chapter ; Yin_MPT ; Yin_icecreamcone ; Yin_g ; Bell_JPC_74_Ising_ferri_1 and is given by
@@ Appendix A The Method
which becomes non-analytic. The difference between Eq. (12) and Eq. (14) takes place in a region of $(T_{0}-\delta T,T_{0}+\delta T)$, where the crossover width $2\delta T$ can be estimated by $|\Upsilon_{-}|=2e^{-2\beta J}\boxed{+-}$ at $T_{0}\pm\delta T$. Following Ref. Yin_MPT, , an alternative and consistent way to measure $\delta T$ is to find such an order parameter that has a well-defined value space of $[-1,1]$ with the value 00 meaning $T_{0}$ and its inverse slope at $T_{0}$ meaning $\delta T$. It is the magnetization of the ordinary spins given by
$\langle\sigma_{i}\rangle=-\frac{\partial f}{h\,\partial\mu_{a}}=\frac{\Upsilon_{-}}{\sqrt{\Upsilon_{-}^{2}+4e^{-4\beta J}\boxed{+-}^{2}}}.$ (15)
@@
Again, it is clear that the crossover width $2\delta T$ decreases exponentially as $J$ increases for fixed finite $T_{0}$. This order parameter provides an accurate, convenient, and microscopic description of CPT. The use of the CPT order parameter accelerates the finding of CPT.
@@
$\begin{split}\delta T=e^{-\frac{2J}{k_{\mathrm{B}}T_{0}}}\frac{2k_{\mathrm{B}}T_{0}^{2}\boxed{+-}}{-\Upsilon_{+}}\biggl{[}h\mu_{a}&+(2J_{1}+h\mu_{b})\tanh(2\beta J_{1}+\beta h\mu_{b})\\ &-(2J_{1}-h\mu_{b})\tanh(2\beta J_{1}-\beta h\mu_{b})\biggr{]}^{-1}_{\beta=\frac{1}{k_{\mathrm{B}}T_{0}}}.\end{split}$ (28)
@@
Appendix B The Magnetization
It is noteworthy that the magnetization $m=\mu_{a}\langle\sigma_{i}\rangle+\mu_{b}\sum_{k}\langle s_{i,k}\rangle$ offers less clear evidence of the CPT at $T_{0}$. It was shown that for $J_{2}/J_{1}=1.95$ and $J=0$, no traces of the pseudo-transition could be found in the density plot of $m$ [see Fig. 8(a)], while the pseudo-transition could be clearly seen in the respective density plots of the entropy, susceptibility, and specific heat 016_Strecka_book_chapter . This insensitivity of $m$ can be explained by the fact that $J_{2}/J_{1}=1.95$ is near the FRI-FRU phase boundary where $m$ is almost the same on both sides with exactly two spins up and one down per unit cell in the ground state as shown in Fig. 2. This feature remains unchanged for $J>0$ [see Fig. 8(b)]. When moving away from the FRI-FRU phase boundary, $m$ becomes a better indicator of the CPT; for example, for $J_{2}=0$, the hidden frustration involves the SPP phase and the CPT can be revealed in the density plot of $m$ except for the weak $h$ region [see Fig. 8(d)]. The comparison between $m$ and $\langle\sigma_{i}\rangle$ the OP will be addressed in more details in subsequent publications Yin_Ising_III .
@@ Appendix B The Magnetization
Figure 8: The $T-h$ phase diagrams from the density plot of the magnetization per spin $m/3=(\mu_{a}\langle\sigma_{i}\rangle+\mu_{b}\langle s_{i,1}+s_{i,2}\rangle)/3$ for (a) $J_{2}/J_{1}=1.95$ and $J=0$, (b) $J_{2}/J_{1}=1.95$ and $J=0.4S^{2}$, (c) $J_{2}/J_{1}=0$ and $J=0$, (d) $J_{2}/J_{1}=0$ and $J=20S^{2}$. $J_{1}=-S^{2}$ and $\mu_{a}=\mu_{b}=S$ where $S=1/2$.